OpenStudy (anonymous):

It took Matt twice the time to drive from home to work as it took him to drive back home from work by the same route.If Matt averaged 30 miles per hour on his drive to work,how was his average speed for the entire round trip? A. 30 mph B. 25 mph C. 40 mph D. 15 mph

OpenStudy (phi):

home to work: 2T hours at 30 mph work to home: T hours at unknown speed use rate*time= distance 2T*30= d d= 60T this is the distance from home to work 2d is the round-trip distance

OpenStudy (phi):

average speed for the entire round trip? average speed is total distance / total time total distance is 2d= 120T total time is 2T+T= 3T average speed is \[ \frac{120T}{3T}= 40 \]